### v1

# Pensieve Grader: An AI-Powered, Ready-to-Use Platform for Effortless Handwritten STEM Grading

Abstract
Grading handwritten, open-ended responses remains a major bottleneck in large university STEM courses. We introduce Pensieve (https://www.pensieve.co), an AI-assisted grading platform that leverages large language models (LLMs) to transcribe and evaluate student work, providing instructors with rubric-aligned scores, transcriptions, and confidence ratings. Unlike prior tools that focus narrowly on specific tasks like transcription or rubric generation, Pensieve supports the entire grading pipeline—from scanned student submissions to final feedback—within a human-in-the-loop interface.
Pensieve has been deployed in real-world courses at over 20 institutions and has graded more than 300,000 student responses. We present system details and empirical results across four core STEM disciplines: Computer Science, Mathematics, Physics, and Chemistry. Our findings show that Pensieve reduces grading time by an average of 65%, while maintaining a 95.4% agreement rate with instructor-assigned grades for high-confidence predictions.
Pensieve Grader: An AI-Powered, Ready-to-Use Platform for Effortless Handwritten STEM Grading
Yoonseok Yangthanks:  Equal Contribution.thanks:  Corresponding author   Minjune Kim11footnotemark: 1   Marlon Rondinelli   Keren Shao Pensieve Inc. {yoon, minjune, marlon, keren}@pensieve.co
1 Introduction
Over the past decade, enrollment in large introductory STEM courses has grown significantly across universities. Courses such as introductory computer science, calculus, and data science routinely enroll hundreds to thousands of students each term. While this scale expands access to foundational knowledge, it also imposes major instructional burdens—chief among them is grading.
Rubric-based grading is a standard approach in post-secondary education, designed to ensure fairness and consistency. Instructors assess student responses against predefined criteria and often refine rubrics iteratively as common error patterns emerge. However, this process is labor-intensive, particularly for handwritten or open-ended responses typical in mathematics, physics, and engineering. Poor handwriting, disorganized reasoning, and irrelevant but superficially correct work often hinder the grading process. These challenges are further exacerbated by increasing class sizes, which delay feedback and overload instructional teams.
Tools like Gradescope Singh et al. (2017) have streamlined parts of the workflow by digitizing submissions and enabling structured rubric application. Yet human graders remain essential for interpreting complex, unstructured student work. Traditional autograders fall short in these settings due to their inability to process free-form inputs or reason about language semantics.
Recent advances in generative AI—especially large language models (LLMs)—offer a promising alternative. LLMs have demonstrated strong capabilities in natural language understanding, zero-shot reasoning, and feedback generation. In domains such as Automated Short Answer Grading (ASAG) and Automated Long Answer Grading (ALAG), LLMs have been used to transcribe student responses, generate rubrics, produce grading rationales, and provide targeted feedback with minimal fine-tuning.
Despite this progress, existing systems often address isolated tasks and do not provide end-to-end support for grading workflows as used in real classrooms. To be practically viable—particularly for instructors accustomed to traditional online grading platforms—an LLM-based grading system must satisfy several requirements:
  1. 1.
    Raw Input Processing: Handle handwritten submissions directly from images or PDFs, without assuming access to clean transcriptions.
  2. 2.
    Rubric Prediction and Refinement: Generate rubric items from problem statements and optionally from reference solutions, and update them based on observed error patterns.
  3. 3.
    Partial Credit Assignment: Select rubric items and assign scores granularly, rather than outputting only final grades or free-form justifications.
In this paper, we present Pensieve (https://www.pensieve.co), an LLM-powered grading system designed to meet these needs. Pensieve has been deployed in real-world classrooms, processing over 300,000 student responses across more than 20 institutions. Our system integrates transcription, rubric induction, and rubric-aligned grading within a familiar, human-in-the-loop interface that aligns with existing grading workflows. We evaluate its performance through empirical studies measuring alignment with instructor grading and reductions in grading time.
Our results show that Pensieve can automate a substantial portion of the grading pipeline while preserving grading accuracy and instructor oversight. This work offers a scalable, practical solution for LLM-assisted grading and represents a step toward closing the gap between AI capabilities and instructional needs in large-scale education.
2 Related Works
We review prior work on rubric-based grading, focusing on systems that leverage large language models (LLMs) across the grading pipeline.
2.1 Image-to-Text Transcription
Accurate transcription of handwritten student work, especially in STEM subjects, remains a major challenge. Traditional OCR systems Smith (2007); Garain and Chaudhuri (2007) often fail to interpret complex mathematical notation due to symbol variety and layout ambiguity. Recent approaches Mok et al. (2024) use multimodal LLMs (e.g., GPT-4V) for image-to-text conversion, achieving strong results on clean inputs but still struggling with subtle distinctions—for example, mistaking $3]{2}$ for $3\sqrt{2}$—which can mislead downstream grading. To mitigate hallucinations, hybrid pipelines (e.g., combining Mathpix and LLMs Kortemeyer et al. (2024)) have been proposed, though they trade off scalability due to longer processing times.
2.2 Rubric Generation and Refinement
In traditional settings, graders iteratively adapt rubrics based on observed error patterns. LLMs offer a promising solution to bootstrap or dynamically revise rubrics. Some systems manually refine rubric items post hoc based on autograding inconsistencies Kortemeyer and Nöhl (2025), while others incorporate human feedback or reinforcement learning to improve rubric clarity and coverage Chu et al. (2024); Li et al. (2025). However, studies show that model-generated rubrics often lack the logical depth and contextual awareness of human-designed ones Wu et al. (2025); Shankar et al. (2024), and iterative rubric self-evaluation by LLMs can introduce bias.
2.3 Rubric Item Selection and Prompt-Based Reasoning
In ALAG, especially for math and physics, assigning partial credit is essential for fair grading Singh et al. (2017). A common approach is to break rubrics into binary items and evaluate each independently. Sonkar et al. Sonkar et al. (2024) follow this method in chemistry grading, improving clarity and consistency. Extensions include ensemble voting and uncertainty flags (e.g., “unsure”) to reflect model confidence Chen and Wan (2025).
Prompt engineering further improves grading quality. Chain-of-thought (CoT) prompts Lee et al. (2024) guide LLMs to reason step by step, while Chen et al. Chen and Wan (2024) introduce scaffolded CoT to explicitly align model reasoning with rubric criteria. Xie et al. Xie et al. (2024) enhance inter-response consistency by prompting over batches of student answers. Retrieval-augmented generation (RAG), as shown by Qiu et al. Qiu et al. (2024), reframes grading as a question-answering task grounded in reference rubrics and solutions.
2.4 Grading Interfaces and User Experience
Despite advances in grading accuracy, many systems lack usable interfaces for non-technical instructors. Tools like BodhiTree-EvalPro Nagakalyani et al. (2025) integrate LLMs but are often limited to structured domains (e.g., code) or rely on unfamiliar UI paradigms.
Pensieve addresses this gap by embedding LLM-assisted grading within a familiar, industry-standard interface. It supports rubric editing, partial credit selection, and instructor review, while providing class-level grading analytics such as mean and standard deviation—making AI-powered grading accessible to real-world classrooms.
Figure 1: Pensieve Grading Interface. On the left, instructors can view the student work and an AI-generated transcription. On the right, instructors can view the AI-generated grade, a confidence level, selected rubrics, AI-generated comments, and AI summary for the reviewer.
3 System Details
This section provides a detailed overview of Pensieve’s design and workflow. The system builds on familiar grading practices while integrating LLM-powered automation to streamline the grading of open-ended student responses.
Instructors can upload student work scans in bulk, or students may upload their assignments individually. When uploaded in bulk, Pensieve automatically matches scans to students by recognizing handwritten names, reducing manual overhead during setup.
For each question, instructors specify the format as one of the following:
  • Single Select Multiple Choice (SSMC)
  • Multi-Select Multiple Choice (MSMC)
  • Drawing
  • Text/Code (all other open-ended formats)
While Pensieve supports autograding for multiple-choice formats, this paper focuses on open-ended responses—such as handwritten solutions, text, or code—where LLMs provide the greatest benefit. We now describe the core AI-assisted components that distinguish Pensieve from traditional grading systems.
3.1 AI Transcription
For open-ended Text/Code questions, Pensieve first transcribes student responses from scanned images. This transcription step uses OCR and language models to convert handwritten or typed text into machine-readable input.
Each transcription is assigned a confidence level—high or low—based on model certainty. Instructors are encouraged to verify low-confidence transcriptions by inspecting the original student work to ensure accuracy before proceeding with grading.
Figure 2: Sample calibration result. Generated grading wisdoms (gray bullet points) supplement the rubric items to help the AI grader interpret instructor expectations more precisely.
3.2 AI Rubric Generation & Calibration
A key differentiator of Pensieve is its ability to generate and refine grading rubrics using LLMs. Each rubric consists of one or more items, with each item assigned a point value and a descriptive label. Rubrics can follow either a subtractive scheme—deducting points for common errors—or an additive scheme—awarding points for correct components. Minimum and maximum total scores can also be specified.
After the initial rubric generation, Pensieve supports a calibration process to fine-tune the rubric and grading behavior based on instructor input. This is particularly useful for open-ended questions, where rubric interpretation can be subjective. For example, instructors may differ in how they interpret vague labels such as “insufficient explanation” (see Figure 2).
To calibrate the AI grader, instructors review a small number of AI-graded examples and correct any mistakes by adjusting the rubric selections. Pensieve then uses these corrections to refine both the rubric interpretation and the grading logic.
During calibration, the system synthesizes grading wisdoms—detailed, explainable instructions derived from discrepancies between the AI’s initial grades and the instructor’s corrections. These wisdoms help the AI capture grading nuances more accurately and are editable by instructors for greater control.
Over time, Pensieve also learns from historical rubric usage and calibration patterns to generate better-aligned rubric suggestions for future assignments, significantly reducing instructor workload and promoting grading consistency.
3.3 AI Grading Confidence
Pensieve assigns each AI-generated grade a confidence level—high, medium, or low. Instructors can use these confidence scores to customize the level of oversight they apply to AI-generated results.
For lower-stakes assignments like homework, instructors may choose to rely entirely on the AI grader. For high-stakes assessments such as exams, instructors can limit trust to high-confidence grades—tuned to match human-level accuracy—while manually reviewing lower-confidence results.
To support this flexible oversight, instructors can configure Pensieve’s behavior based on confidence levels:
  • Transcription visibility: Show or hide the AI-generated transcription of student responses.
  • Autograde visibility: Show or hide the AI-assigned grade.
  • Automatic review: Require manual confirmation of AI-assigned grades.
3.4 AI Feedback
In large STEM courses, students often receive limited feedback on open-ended questions due to the grading burden. Pensieve addresses this by allowing instructors to generate individualized feedback after grading.
Once grading is complete, the system can generate comments that explain the student’s mistakes, referencing the associated rubric items. Instructors can also provide custom prompts to guide the style, tone, or content of the generated feedback.
3.5 AI Summary
To improve grading transparency and help instructors understand AI decisions, Pensieve generates a concise summary of each student response. These summaries highlight the key reasoning behind the selected rubric items and flag major errors, enabling instructors to verify the AI’s work more efficiently.
4 Evaluation
We evaluate the system from two key perspectives: (1) the growth in the number of autograding tasks completed using Pensieve, and (2) the amount of grading time saved for instructors.
Figure 3: Cumulative Number of Questions Graded Using Pensieve (Jan–Jun 2025)
4.1 Usage Patterns and Subject-wise Distribution
Figure 3 shows a clear upward trend in the number of autograded questions across the Winter and Spring quarters (or Spring semester). Usage spikes align with midterm and final exam periods, indicating that instructors primarily rely on the system during high-stakes assessments—likely to reduce turnaround time and grading workload while accelerating student feedback.
Figure 4: Distribution of subjects that used Pensieve
Figure 4 illustrates the subject-wise breakdown of autograded questions. Computer Science dominates, as expected, due to the structured nature of code-based answers and the minimal need for handwriting transcription. More notably, Mathematics and Physics also account for a substantial portion of usage, demonstrating that Pensieve’s LLM-powered transcription and rubric reasoning pipelines can effectively handle symbol-heavy, multi-step problems at the university level. As discussed earlier, these subjects present significant challenges due to ambiguous notation and complex contextual reasoning, yet the system still meaningfully reduces grading effort.
Because this analysis focuses solely on free-form responses, subjects that typically rely on multiple-choice or short-answer formats (e.g., Chemistry, Biology) appear underrepresented. However, the system remains applicable to mixed-format assessments that include even a small number of open-ended questions.
Figure 5: Average Grading Time per Submission: With AI vs. Without AI
Overall CS Math Phys Chem
Acc(%) 95.4 95.8 93.5 94.5 97.5
Table 1: Accuracy of High-Confidence AI Grades.
4.2 Time Savings Across Subjects
Figure 5 presents the estimated number of grading hours saved per assignment across different subjects. Although the exact savings depend on factors such as assignment length, class size, and subject complexity, the reductions are consistently significant. In high-enrollment courses, the benefits are particularly striking—often saving instructors dozens of hours per assignment.
We estimate the time savings using a straightforward model, where the total grading time without automation is approximated as follows:
$t_{\text{without autograder}}=t_{avg}IJ$
where $t_{avg}$ is the average grading time per question (based on user-reported or observed metrics), $I$ is the number of students, and $J$ is the number of problems. When using the autograder, instructors are expected to (1) review all low-confidence outputs and (2) verify a small portion of high-confidence outputs that may still contain errors—typically around 5%, as indicated in Table 1. Thus, the adjusted grading time becomes:
$t_{\text{with autograder}}=t_{avg}(IJ-c)$
where $c$ is the number of correctly autograded, high-confidence responses.
Across real-world assignments, this results in time savings between 40% and 80%, depending on the quality of the rubric, the clarity of student responses, and the domain. When scaled to classes with 200–300 students, this reduction not only saves labor but also accelerates feedback cycles—benefiting student learning and improving instructional agility.
Conclusion
We presented a system that leverages AI to significantly reduce grading time—by an average of 65%—across more than 300,000 open-ended questions from four major college-level STEM domains: Computer Science, Mathematics, Physics, and Chemistry. This reduction is achieved without any meaningful loss in grading accuracy, even for complex, free-form responses.
Pensieve integrates large language models into every stage of the grading workflow, from rubric generation and calibration to transcription, grading, and feedback. Its adaptive learning and instructor-in-the-loop calibration mechanisms ensure that grading aligns with diverse teaching styles while minimizing manual effort.
Designed for practical deployment, Pensieve is readily usable by any instructor. By automating repetitive evaluation tasks, our system enables educators to focus more on high-impact activities such as student interaction and support—while also providing faster, more detailed feedback to learners.